OpenStudy (anonymous):

Convert to a fraction in lowest terms. 33.4%

hero (hero):

33.4/100 = 334/1000 = 167/500

OpenStudy (anonymous):

I have a question for you, how did you get from 33.4/100 to 33.4/1000? I really need to learn how to do this and I have to show my work. Thank you for the fast reply!

hero (hero):

Just move the decimal place over on top and bottom

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):